# Potentials on the conformally compactified Minkowski spacetime and their application to quark deconfinement

### v1

@@ Abstract
We study a class of conformal metric deformations in the quasi-radial coordinate parameterizing the 3-sphere in the conformally compactified Minkowski spacetime $S^{1}\times S^{3}$. After reduction of the associated Laplace-Beltrami operators to a Schrödinger form, a corresponding class of exactly solvable potentials (each one containing a scalar and a gradient term) is found. In particular, the scalar piece of these potentials can be exactly or quasi-exactly solvable, and among them we find the finite range confining trigonometric potentials of Pöschl-Teller, Scarf and Rosen-Morse. As an application of the results developed in the paper, the large compactification radius limit of the interaction described by some of these potentials is studied, and this regime is shown to be relevant to a quantum mechanical quark deconfinement mechanism.
@@ 1 Introduction
Our main goal in the present work is to construct, through arbitrary $M^{c}$ metric deformations, a new family of scalar potentials whose exact ground state solutions are determined by the corresponding conformal Laplacians. Among them one finds the already mentioned trigonometric Rosen-Morse potential, the (well known from super-symmetric quantum mechanics) exactly solvable trigonometric Pöschl-Teller and Scarf potentials, and other potentials which are only quasi-exactly solvable. We refer to all of them as ‘induced’ or ‘mean field’ scalar potentials. Furthermore, we explore importance of the compactification radius in the physics of confinement, for the particular case of the trigonometric Rosen-Morse potential interpreted as a color-electric charge dipole potential. We show that for large compactification radii, as they can occur at high temperatures in heavy ion collisions, the color–electric charge dipole potential can collapse to one of Coulomb-type, thus locally enabling the formation of multi-$q\bar{q}$ systems of Rydberg-atom type. In this way, at some large compactification radii, the Rydberg quark, occupying an orbital with very high principal quantum number, can become loosely bound and be knocked out in an inelastic scattering event, thus becoming accessible to observation. Therefore, we address from first principles the fundamental question of the possible existence of a deconfined phase at very high temperatures.
@@ 5 Application to quark deconfinement
[…]
@@ 5.3 Application to hadron spectra
[…]
@@ 5.4 The color electric charge dipole potential at asymptotically large compactification radii
$\overline{u}$$d$$b$$\overline{b}$$\overline{u}$$b$
@@ 6 Conclusions and outlook
  • This process allows for the formation of hydrogen-like quarkish Rydberg-states, known to be easily ionizable. Indeed, the Rydberg quark, occupying an orbit with a high principal quantum number, behaves approximately as a free particle and can be knocked out in an inelastic scattering event, thus becoming observable.

### v2

@@ Abstract
We study a class of conformal metric deformations in the quasi-radial coordinate parameterizing the 3-sphere in the conformally compactified Minkowski spacetime $S^{1}\times S^{3}$. Prior to reduction of the associated Laplace-Beltrami operators to a Schrödinger form, a corresponding class of exactly solvable potentials (each one containing a scalar and a gradient term) is found. In particular, the scalar piece of these potentials can be exactly or quasi-exactly solvable, and among them we find the finite range confining trigonometric potentials of Pöschl-Teller, Scarf and Rosen-Morse. As an application of the results developed in the paper, the large compactification radius limit of the interaction described by some of these potentials is studied, and this regime is shown to be relevant to a quantum mechanical quark deconfinement mechanism.
@@ 1 Introduction
Our main goal in the present work is to construct, through arbitrary $M^{c}$ metric deformations, a new family of scalar potentials whose exact ground state solutions are determined by the corresponding conformal Laplacians. Among them one finds the already mentioned trigonometric Rosen-Morse potential, the (well known from super-symmetric quantum mechanics) exactly solvable trigonometric Pöschl-Teller and Scarf potentials, and other potentials which are only quasi-exactly solvable. We refer to all of them as ‘induced’ or ‘mean field’ scalar potentials. Furthermore, we explore importance of the compactification radius in the physics of confinement, for the particular case of the trigonometric Rosen-Morse potential interpreted as a color-electric charge dipole potential. We show that for large compactification radii, as they can occur at high temperatures, the color–electric charge dipole potential can collapse to one of Coulomb-type, thus locally enabling the formation of multi-$q\bar{q}$ systems of Rydberg-atom type. In this way, at some large compactification radii, the Rydberg quark, occupying an orbital with very high principal quantum number, can become loosely bound and be knocked out in an inelastic scattering event, thus becoming accessible to observation. Therefore, we address from first principles the fundamental question of the possible existence of a deconfined phase at very high temperatures.
@@ 4.3.3 The trigonometric Scarf potential
4.3.4 The potential of the MIC-Kepler problem [^1] This result has been obtained after submission of galley proofs, so it does not appear in the published version.
Combining the conformal metric deformation for the trigonometric Rosen-Morse with the one for the Pöschl -Teller potential (though with $\sin\chi$ in place of $\cos\chi$ in the latter) i.e. by $f(\chi)=\beta/2\ln\sin\chi+\alpha_{K}\chi/2$, we obtain
$V_{MIC-K}(\chi)=\frac{\ell(\ell+1)}{\sin^{2}\chi}+\frac{\mu^{2}}{\sin^{2}\chi}+\alpha_{K}(K+1)\cot\chi-\frac{\beta^{2}}{4}+\frac{\alpha_{K}^{2}}{4}-\beta(K+1)-(K+1)^{2},$ (79)
with $\beta$ constant and $\mu^{2}=\beta/2(\beta/2-1)+\beta(K+1)$. This is the exactly solvable MIC-Kepler potential describing on $S^{3}$, according to [18], the interaction of Schwinger’s dyons [19], particles equipped by both electric and magnetic monopole charges, in which case $\mu$ is integer or semi-integer. The present consensus is that quarks can not be dyons because the electric charge of a dyon has to be integer, versus semi-integer in quarks. Our model shows that there exists a metric on the compactified Minkowski spacetime which can give rise to a mean field potential of the dyon–dyon type. As to what extent such a metric could be relevant to physics, this can be decided only upon comparison of its predicted hadron magnetic form factors to data.
@@ 5 Application to quark deconfinement
5.1 The conformal wave equation with the one dimensional color-electric charge dipole potential in units of MeV2
@@ 5.3 Application to hadron spectra
The values of their positions have then be adjusted by adding to (88) the constant term, $C(R)=1.1873$ GeV2.
@@ 5.4 The color electric charge dipole potential at asymptotically large compactification radii
$\overline{u}$$d$$b$$\overline{b}$
$\overline{u}$$b$
@@ 6 Conclusions and outlook
  • This process allows for the formation of hydrogen-like quarkish Rydberg-states, known to be easily ionizable. Indeed, the Rydberg quark, occupying an orbit with a high principal quantum number, behaves approximately as a free particle and can be knocked out in an inelastic scattering event.